### v1

# Self-Controlled Diffusion for Denoising in Scientific Imaging

Abstract
This paper presents a novel approach for denoising Electron Backscatter Diffraction (EBSD) patterns using diffusion models. We propose a two-stage training process with a UNet-based architecture, incorporating an auxiliary regression head to predict the quality of the experimental pattern and assess the progress of the denoising process. The model uses an adaptive denoising strategy, which integrates quality prediction and feedback-driven iterative denoising process control. This adaptive feedback loop allows the model to adjust its schedule, providing fine control over the denoising process. Furthermore, our model can identify samples where no meaningful signal is present, thereby reducing the risk of hallucinations. We demonstrate the successful application of diffusion models to EBSD pattern denoising using a custom-collected dataset of EBSD patterns, their corresponding master patterns, and quality values.
1 Introduction
EBSD, or Electron Backscatter Diffraction, is a technique used in electron microscopy to study the crystallographic properties of materials at the nanoscale. In this method, electrons from an electron beam are diffracted by crystallographic planes, creating a diffraction pattern that can be analyzed to determine the crystal orientation. Due to its high resolution, EBSD can provide detailed information about grains, including those that are nanometer-sized.
Figure 1: a) Pattern formation as a tangential projection of diffraction sphere on a detector plane; b) Inverse Polar Figure (IPF) map visualising the orientation of a grain by the color of the crystal face pointing perpendicular to the image plane. Each point of IPF map is the result of EBSD pattern indexing.
EBSD patterns appear as grayscale images featuring diffraction lines, known as Kikuchi lines. Each EBSD pattern represents a portion of the entire diffraction sphere. Similar to maritime navigation, where specific sections of the sky help determine one’s location, the Kikuchi pattern is utilized to identify the crystallographic orientation of individual crystallites (Figure 1). Process of determination of crystallographic orientation by analysis of Kikuchi patterns is called indexing (assigning crystallographic indexes to every Kikuchi line appear at the pattern).
The quality of EBSD results depends on several parameters, including those related to the electron microscope (such as probe current, accelerating voltage, working distance, and magnification), the type of detector (e.g., CCD or single electron detectors), and, most importantly, the sample itself (surface roughness, composition, crystallization degree, chemical and thermal stability and so on) [18, 7]. Stable and well-crystallized inorganic materials, like steel and its alloys or alumina, produce clear and distinguishable Kikuchi patterns [4, 3]. In contrast, some beam-sensitive materials, such as perovskite which become now the most promising materials for the green energy transition as active layer material in solar cells, or certain metals in thin film form, generate weak and noisy patterns [16, 7].
The analysis of weak patterns is a complex task. Currently, there are three main approaches to patterns indexing: Hough indexing (HI) [8, 15], Dictionary Indexing (DI) [20] and relatively recently developed Spherical indexing (SI) [17].
Hough Indexing converts an EBSD pattern into Hough space to summarise the intensity of the entire Kikuchi line into a single point. This representation allows for the detection of that line in the original pattern. By analysing the overall position of all lines, the crystal orientation can be determined. However, a major drawback of this method is its requirement for high-quality patterns, which limits its applicability to beam-sensitive materials.
The Dictionary Indexing (DI), on the other hand, offers much greater precision in determining orientation, even for somewhat noisy patterns. This method involves using an algorithm that compares complete experimental EBSD patterns against simulated patterns (Figure 2) [20]. Despite its advantages, the DI is quite slow because it necessitates the generation of a large number of parts of the simulated Master Pattern and massive amount of calculations during analysis.
Currently, the most advanced indexing method is the Spherical Indexing (SI). Similar to the DI, it requires the generation of a complete diffraction picture known as the Master Pattern. However, instead of calculating a dot product for each experimental pattern and piece of the Master Pattern, the Spherical Indexing utilizes spherical functions, resulting in higher precision and much faster analysis. However, it still necessitates considerable processing power.
Figure 2: An example of Master Pattern (MP) and its components identified during Spherical Indexing. The same components of MP served as ground truth in model training.
However, like all indexing techniques, SI’s effectiveness relies heavily on the pattern preprocessing procedure. Adjustments to the image processing steps can lead to significant changes in results, and it is not always clear which procedures should be applied and in what order. Presently, achieving optimal results often depends on experience and, at times, luck, as the signal quality can vary from sample to sample and from measurement to measurement. Additionally, the current capabilities of SI allow for the indexing of highly noisy patterns, where not a single trace of a Kikuchi line is detectable by the naked eye. Therefore, it can be difficult to determine whether the indexing is truly accurate.
In this context, EBSD pattern restoration might be a valuable part of the EBSD post-analysis. The ability to restore noisy patterns opens up possibilities for improved analysis in three key ways:
  • enhance pattern processing to enable indexing of patterns that were previously unindexable;
  • reduce exposure time in situations where prolonged signal collection might damage samples, all while maintaining quality;
  • make indexing more reliable, consistent, and less dependent on the preprocessing of patterns.
The last two points can be partially addressed by utilizing advanced detectors, such as single-particle-counting detectors (e.g., Medi/Timepix series), which provide noiseless digital integration (counting) and high detection efficiency [11]. A particle-counting detector functions such that each pixel serves as a detector. Each photon (or electron) that reaches a pixel is compared to a predefined threshold, and a digital counter is incremented accordingly. These detectors are particularly effective when signal intensity is low and collecting substantial signal for adequate statistical noise suppression would damage the sample. Despite their advantages, they are often insufficient for beam-sensitive materials, and the resulting images still require further processing before they can be indexed. Though, combination of them, such as single photon detector and denoising system restoring noisy patterns to a more clear EBSD patterns, can become a very powerful solution helping to reveal new insights about materials and making EBSD method more common and easy in use making it more accessible for more researchers.
Computer vision approaches offer a promising way to denoise EBSD patterns and extract more valuable information from them. Currently, several methods are employed for this purpose: Autoencoders, Generative Adversarial Networks (GANs), and diffusion-based learning techniques. While researchers have applied the first two approaches to denoise EBSD patterns, their effectiveness is constrained by pattern quality, which must remain sufficiently high to ensure stable results and prevent hallucinations during restoration. Diffusion learning offers an advantage through its step-by-step procedural approach. However, as we will demonstrate in the Related Works section, even these sophisticated techniques can produce hallucinations, particularly when processing content-simple images like EBSD patterns.
To address these challenges and improve the reliability of EBSD pattern denoising, this paper aims to develop a control mechanism that limits the restoration process while preserving vital information in grayscale, line-based EBSD images. Our contributions include:
  • We introduce a feedback-driven, adaptive denoising approach that integrates quality prediction to control the iterative process. To our knowledge, this is the first successful application of diffusion models to EBSD pattern denoising.
  • Our model incorporates a mechanism for distinguishing between meaningful signal and noise, helping to flag potential hallucinations.
2 Related Works
Improving the quality of collected EBSD patterns (EBSP) using Deep Learning (DL) approaches is an active topic in recent years. A review of existing literature reveals two main strategies of doing this:
  • Directly enhancing EBSP to improve the quality of standard indexing procedures. This involves collecting EBSP with sequentially improved visibility of Kikuchi lines or an enhanced Signal-to-Noise Ratio (SNR). These DL optimized patterns can then be analyzed using SI/DI/HI for orientation determination;
  • Indexing by SI, DI, or HI methods EBSP that have been processed using traditional imaging techniques. This step is followed by the enhancement of the resulting Inverse Pole Figure (IPF) maps through Deep Learning.
The first approach appears to be the more convenient choice for enhancing the quality of EBSD results. Since the pattern dataset serves as the foundation for the entire process, starting with this step could lead to significant overall improvements in the analysis.
The initial studies on denoising EBSD patterns utilized UNet-based autoencoders [2] where the authors achieved a commendable quality of restored patterns. At that time, Spherical Indexing was not widely adopted, and the initial patterns they worked with were often poorly indexed using Hough Indexing. Nowadays, SI can effortlessly index patterns of such quality. Consequently, a significant portion of this work is focused on enhancing the precision of strain analysis, which requires much higher-quality patterns.
It’s important to highlight that there are very few papers dedicated to pattern restoration when the patterns are predominantly noise and nearly devoid of detectable lines to the human eye. One notable exception is the work by [12], where the patterns are nearly lost, although an experienced user can still identify some Kikuchi lines. The lack of research addressing such low-quality patterns arises from the lack of methods to verify whether predictions are accurate and the challenges associated with collecting ground truth data.
The second approach is less common and requires more effort in data collection, as an Inverse Pole Figure (IPF) map consists of thousands of EBSD patterns. To obtain reliable results, a large number of experiments must be conducted to create a sufficiently sized dataset based on IPF maps. This process often involves significant data augmentation, which can lead to the sacrifice of some details within the data. In the work by [14], the IPF maps were restored using the Diffusion approach. To stabilize the diffusion model, they utilized a Reduced-Order Generalized Spherical Harmonic (ROGSH), which incorporates additional information to enhance restoration performance.
Diffusion models have become one of the most powerful approaches for image restoration, particularly for denoising, due to their ability to restore images gradually and more carefully compared to alternatives such as GANs or VAEs. However, this stability comes at the cost of significant computational resources needed for the restoration process that might be critical in terms of using such networks for ”every-day” EBSD dataset restoration what usually contains thousands or even tens of thousand images.
Nevertheless, the main advancement of diffusion models that enhance image restoration precision is their approach to handling noisy images. Unlike classical models such as GANs or VAEs, which aim to transform a noisy sample directly into a denoised one, diffusion models transform a noisy image into a less noisy version in several steps which is especially valuable in application to scientific images which are formed due to physical properties of samples. This gradual restoration occurs within the space of images, transitioning from a noisy state to a denoised one. This approach was originally proposed not just for diffusion models, but as a more general method for restoring images without the need to train separate models for every restoration problem [10].
The diffusion models are represented by several types of approaches. The Denoising Diffusion Probabilistic Model introduced by [5] uses Markov chains to model the forward and reverse process. Step by step it applies noise (usually Gaussian) during forward process. Then in backward process the model is learning noise distribution parameters such as mean $\mu_{\theta}(x_{t},t)$ and variance $\beta_{\theta}(x_{t},t)$ to be able to restore image from noisy state back to original state.
Despite their popularity in many imaging tasks recently, the diffusion models are still not widely used for scientific image and especially EBSD images denoising. However, it proved its efficiency in denoising regular images like photographs [19]. The crucial difference in the [19] work is in method of sampling they applied. In previous works, every next image in the denoising pipeline is a combination of restored image from the previous state and normalized noise which is left after subtraction of restored image from the noisy one (1). In the [19] work, every next image is an image from the previous step and subtracted normalized difference between original noisy image and a restored image from the previous step (2). Changing the sampling method has helped stabilize the image restoration process.
$x_{t-1}=S_{\theta}(x_{t},t)+\frac{(t-1)}{T}\cdot[x_{T}-S_{\theta}(x_{t},t)]$ (1)
$x_{t-1}=x_{t}-\frac{1}{T}\cdot[x_{T}-S_{\theta}(x_{t},t)]$ (2)
As mentioned above, the main challenge for all denoising approaches is ensuring the stability of the image restoration process. In [9] was demonstrated, that the restored image is not a single fixed image but rather a variety of potential images, forming a data manifold. A promising approach suggested by [1] involves using an adaptive level of noise during each step of the reverse process, rather than applying a constant noise level. This adaptive noise level is estimated at each step based on the distance of the restored image $x_{t}$ from the noiseless image $x_{0}$. To facilitate this, a separate network has been introduced to learn the level of noise correction, which is defined as follows: $\hat{\sigma}_{t}:=\sigma_{t}[1+\hat{r}_{t}]\approx\text{dist}_{\kappa}(x_{t})/\sqrt{n}$.
3 Dataset
There are EBSD datasets available on the Internet (https://zenodo.org/communities/ebsd/ ). However, these datasets were not used in this work as our approach relies on Master Patterns as a ground truth and employs a patterns quality metric, which is not available in the public data. Therefore, we collected new datasets and generated corresponding pieces of Master Patterns using Spherical Indexing.
To collect the experimental dataset for model training, we utilized a Scanning Electron Microscope Verios 460 from ThermoFisher, equipped with an EDAX Clarity detector utilizing a 4-chip Timepix detector from Amsterdam Scientific Instruments (ASI). The data patterns were gathered using EDAX APEX software with beam currents ranging from 50 to 200 pA and accelerating voltages between 5 and 20 kV. Pixel integration times varied from 30 to 250 ms, with different binning options applied (1, 2, 4, and 5 pixels). We analyzed multiple samples (Nickel ($Ni$), Ruthenium ($Ru$), Gold ($Au$), Silver ($Ag$)) to apply a diverse set of measurement conditions and to include various crystallographic variables, including different crystal lattice parameters and crystal systems (Cubic, Orthorhombic, Hexagonal). In total, the number of experimental patterns used for training was over 163000, with further 16500 patterns used for validation and over 6000 used as a hold-out test set.
3.1 Master patterns
Master patterns for each crystal structure were generated using EDAX OIM 9 software. This software was also utilized to index experimental EBSD patterns using Spherical Indexing. The master patterns identified through this process were exported and used as ground truth for training purposes.
4 Methods
The data in this task differs from traditional image denoising because it includes both ideal images (master patterns) and experimental images with an associated ’confidence index’ that quantifies the reliability of the experimental data. Additionally, within each grain, the computed master patterns are nearly identical, with only minor variations due to defects or crystal lattice bending. This allows experimental patterns within a grain to be considered noisy samples from the same distribution.
4.1 Proposed approach
Our approach combines a diffusion-based denoising model with an integrated quality assessment feedback mechanism. This dual-objective architecture allows the model to simultaneously denoise EBSD patterns while evaluating the denoisability and quality of the input patterns. Critically, our model includes a mechanism to identify and flag patterns that represent pure noise rather than noisy observations of actual crystal structures, thus preventing hallucination artifacts.
(a) Stage 1: Training the denoising neural network. The example illustrates the process for $t=128$.
(b) Stage 2: Training the auxiliary head to predict both the original image quality and denoising progress $r_{denoise}$. Instead of applying an external noise scheduler (as in Stage 1), the pretrained denoising network is used to provide partially denoised images. The function $g_{\theta}$ is trained using a random value $t_{rand}$ to decouple denoising stage prediction from step number embedding.
Figure 3: Proposed two-stage training procedure for image denoising and quality assessment.
4.2 Training procedure
Model training was performed in two stages.
Denoising model. On the first stage we training the main UNet-style encoder-decoder denoising neural network to perform image denoising. The pseudocode for this training procedure is presented in Algorithm 1, which is depicted in Figure 3(a). During training, the network receives a noisy observation $x_{t}$ as input and attempts to predict the corresponding clean image $x_{0}$. The denoising process is guided by a quality score $q$ and a noise scale $s_{\text{noise}}$, with additional Gaussian noise injected at each diffusion step to simulate varying levels of corruption.
The noise injection process follows a diffusion model, where the noisy observation $x_{t}$ is computed as a weighted combination of the clean image $x_{0}$, the noisy input $x$, and an additional noise term $\omega_{t}$ drawn from a standard Gaussian distribution $\mathcal{N}(0,I)$, Eq. 1.
$x_{t}=(1-\alpha)x_{0}+\alpha x+\beta\omega_{t}$ (1)
The weight $\alpha$ is determined by the current diffusion step $t$, relative to the total number of diffusion steps $T$, as $\alpha=\frac{t}{t_{x}}$, where $t_{x}=T\cdot(1-q)$. The noise scale $\beta$ is determined by the factor $m\cdot\frac{s_{\text{noise}}\cdot t}{T}$, where $m$ is drawn from a Bernoulli distribution with probability $p_{\text{noise}}$, indicating whether additional noise should be applied, while $s_{\text{noise}}$ controls the level of additional image degradation.
Notably, the algorithm allows the generation of images that can be more degraded than the observed image $x$, since for each experimental image $t_{x}\leq T$ and the current step $t$ can exceed $t_{x}$, leading to further simulated corruption of the image. This, together with the extra noise enables the model to recover clean images from increasingly corrupted data.
On each step, the network has access to the interpolated noisy image of the step $x_{t}$ and is guided by the original experimental noisy image $x$, which gives the network more context (both are stacked as a two channel image), and the current step $t$. The network’s parameters $\theta$ are updated through backpropagation to minimize the loss function $\mathcal{L}(\hat{x}_{0},x_{0})$, which measures the difference between the predicted clean image $\hat{x}_{0}$ and the true clean image $x_{0}$.
Algorithm 1 Denoising Model Training
1:Encoder-decoder neural network $f_{\theta}$, dataset $\mathcal{D}=\{(x_{0}^{(i)},x^{(i)},q_{i})\}_{i=1}^{N}$, where $x_{0}^{(i)}$ is the clean image (Master Pattern), $x^{(i)}$ is the noisy observation, and $q_{i}$ is the quality of $x^{(i)}$; total number of diffusion steps $T$, noise probability $p_{\text{noise}}$, and noise scale $s_{\text{noise}}$.
2:while not converged do
3:     Sample $(x_{0},x,q)$ from $\mathcal{D}$
4:     Calculate diffusion step of $x$: $t_{x}=T\cdot(1-q)$
5:     Compute interpolation $x_{t}=(1-\alpha)x_{0}+\alpha x$, where $t\sim\mathcal{U}(0,T)$, $\alpha=\frac{t}{t_{x}}$,
6:     Generate additional noise: $\omega_{t}\sim\mathcal{N}(0,I)$
7:     Add noise: $x_{t}=x_{t}+\beta\omega_{t}$ where $\beta=m\cdot\frac{s_{\text{noise}}\cdot t}{T},\quad m\sim\text{Bern}(p_{noise})$.
8:     Normalize: $x_{t}\leftarrow\frac{x_{t}-\mathbb{E}[x_{t}]}{\text{std}(x_{t})}$
9:     Predict the clean image: $\hat{x}_{0|t}\leftarrow f_{\theta}(x_{t},x,t)$
10:     Update parameters of $f_{\theta}$ to minimize loss: $\mathcal{L}(\hat{x}_{0|t},x_{0})$
11:end while
Quality Values Regression. In the second stage of training (Algorithm 2, depicted in Figure 3(b)), we perform quality regression using the trained denoising network in a frozen state. Specifically, the bottleneck feature map, which captures an internal representation of the input tensor, is passed through an auxiliary head for the two-values regression task. The goal is to predict two key values: the original image quality $\hat{q}$, and the progress of the denoising process $\hat{r}_{\text{denoise}}$.
The first predicted value, $\hat{q}$, represents the estimated quality of the original image at the current denoising step, corresponding to the same ground-truth value $q$ from the dataset for all diffusion steps. Since the experimental noisy image $x$ is part of the input tensor at each step, the information necessary for predicting this objective value is already present in the input.
The second predicted value, $\hat{r}_{\text{denoise}}$, estimates the progress of the denoising process. This value depends on the current diffusion step $t$ and provides a measure of how far along the diffusion process is in recovering the clean image. More specifically, $\hat{r}_{\text{denoise}}$ estimates how much noise has been removed by the current denoising step, or in other words, it can be viewed as a measure of how close the image $x_{t}$ is to the final clean state.
The target value for $\hat{r}_{\text{denoise}}$ is constructed using the following rule:
$r_{\text{denoise}}=\begin{cases}1-\frac{t}{T},&\text{if }x\text{ is not noise},\\ 0,&\text{otherwise.}\end{cases}$ (2)
When the image is not fully corrupted (i.e., $x$ is not pure noise), the value is inversely proportional to the diffusion step, with lower $t$ values indicating a more advanced stage of denoising. For fully noisy images, the progress is set to $r_{\text{denoise}}=0.0$. The sample $x$ can represent pure noise in several cases: the dataset includes intentionally collected pure noise EBSD patterns, where noise collected as instrumental noise without a sample material installed, or synthetic noise samples generated during training by augmenting the data with pure noise drawn from a standard Gaussian distribution $\mathcal{N}(0,I)$. For all such cases, both regression targets are set to 0.
The training of the auxiliary head involves direct estimation of the partially denoised image at diffusion step $t$:
$x_{t}=(1-\frac{t}{t_{x}})f_{\theta}(x,x,t_{x})+\frac{t}{t_{x}}x.$ (3)
This component differs from the procedure in Algorithm 1, as the intermediate image state is generated using the actual denoising model rather than a linear interpolation scheduler, utilising the Master Patterns. Since the encoder’s input includes the diffusion step $t$, and since $\hat{r}_{\text{denoise}}$ is a function of $t$, it is crucial to decouple the image generation process during training from the estimation of $\hat{r}_{\text{denoise}}$. This is done as follows: first, $t\sim\mathcal{U}(0,T)$ is used to construct $x_{t}$ and the target value for $\hat{r}_{\text{denoise}}$. Then during the forward pass, which updates the weights of $g_{\theta}$, a different random step $t_{\text{rand}}\sim\mathcal{U}(0,T)$ is used as input to ensure that the auxiliary head relies on the actual visual features rather than the information from $t$-embedding, allowing the value to be used for denoising progress assessment.
As will be shown in the following sections, this two-values regression enables the model to estimate both the initial step at which the denoising process should begin and provide a feedback on the progress of the denoising process.
Algorithm 2 Quality Regression Training
1:Frozen encoder($f_{\theta}^{enc}$)-decoder($f_{\theta}^{dec}$) denoising neural network $f_{\theta}$, auxiliary quality prediction head $g_{\theta}$, dataset $\mathcal{D}$, total number of diffusion steps $T$.
2:while not converged do
3:     Sample $(x,q)$ from $\mathcal{D}$
4:     Calculate diffusion step of $x$: $t_{x}=T\cdot(1-q)$
5:     Get partially denoised image: $x_{t}=(1-\alpha)f_{\theta}(x,x,t_{x})+\alpha x$, where $\alpha=\frac{t}{t_{x}},t\sim\mathcal{U}(0,T)$
6:     Set the target $r_{\text{denoise}}$ as per Eq. 2, using the same $t$ as used for $x_{t}$.
7:     Predict regression values: $(\hat{q},\hat{r}_{\text{denoise}})\leftarrow g_{\theta}(f_{\theta}^{enc}(x_{t},x,t_{\text{rand}}))$, where $t_{\text{rand}}\sim\mathcal{U}(0,T)$ random diffusion step.
8:     Update parameters $g_{\theta}$ to minimize loss: $\mathcal{L}((\hat{q},\hat{r}_{\text{denoise}}),(q,r_{\text{denoise}}))$
9:end while
4.3 Prediction
Conceptually, the prediction Algorithm 3, depicted in Figure 4, is similar to the improved sampling algorithm from [19], with two key novel features arising from the use of the two regression values. First, we estimate the initial diffusion step based on a dry-run of the neural network. In this step, the input timestep is set to $T$ the total number of timesteps used during training.
Second, in contrast to dynamic methods like [1], which adjust the amount of noise removed in each step based on a monotonous denoising process with predicted noise level, we adaptively select the next diffusion step during denoising based on the predicted denoising progress $\hat{r}_{t_{x}}$. This adaptive approach allows our algorithm to dynamically adjust the denoising steps: it can repeat a step if further denoising is required or skip steps if the progress indicates sufficient improvement. To prevent infinite loops, the total number of denoising steps is capped at $T$, and the predicted value $\hat{r}_{t_{x}}$ provides a stopping criterion for the process, allowing it to exit at a desired denoising level $R$ when required by the application.
Moreover, given the definition of $\hat{r}_{t_{x}}$, its value approaching zero can serve as a useful indicator to flag potential hallucinations or unreliable predictions. This is particularly valuable in ensuring the robustness of the denoising process and detecting situations where the model might be overfitting or making implausible predictions. Usage of $\hat{r}_{t_{x}}$ allows getting the best denoised image, based on the evaluation, rather then conventionally selected image from the last denoising step.
Algorithm 3 Denoising Prediction with Feedback
1:Neural network $E_{\theta}$, consists of encoder($f_{\theta}^{enc}$)-decoder($f_{\theta}^{dec}$) denoising neural network, auxiliary quality prediction head $g_{\theta}$, total number of diffusion steps $T$, desired final denoising $R\in[0,1]$
2:$(\hat{q}_{\text{init}},\hat{r})\leftarrow g_{\theta}(f_{\theta}^{enc}(x,x,T))$
3:Start diffusion step $t_{x}=T\cdot(1-\hat{q}_{\text{init}})$; $x_{t}=x$
4:Step counter $n\leftarrow 0$
5:while $n<T$ and $t_{x}>0$ and $\hat{r}_{t_{x}}<R$ do
6:     $(\hat{x}_{0|t_{x}},\hat{q}_{t_{x}},\hat{r}_{t_{x}})\leftarrow E_{\theta}(x_{t},x,t_{x})$
7:     $\alpha_{t}=t_{x}/T$; $\alpha_{t-1}=(t_{x}-1)/T$
8:     $\widetilde{x}_{t}=(1-\alpha_{t})\hat{x}_{0|t_{x}}+\alpha_{t}x$
9:     $\widetilde{x}_{t-1}=(1-\alpha_{t-1})\hat{x}_{0|t_{x}}+\alpha_{t-1}x$
10:     Denoising: $x_{t}=x_{t}-\widetilde{x}_{t}+\widetilde{x}_{t-1}=x_{t}+\frac{1}{T}(\hat{x}_{0|t_{x}}-x)$
11:     Next step as predicted feedback: $t_{x}=T\cdot(1-\hat{r}_{t_{x}})$
12:     Increment step counter $n=n+1$
13:end while
14:Final $\hat{x}_{0}$ selected as $x_{t}$ with the highest $\hat{r}_{t_{x}}$.
Note that the proposed prediction schedule has no reliance on clean images (Master Patterns), which is crucial for the application.
Figure 4: Diagram of the proposed denoising algorithm with feedback.
4.4 Architecture and Implementation details
All experiments used $128\times 128$ images and a standard UNet architecture with two residual layers per block, following [5, 19]. A key deviation was downscaling only in the first encoder block. The auxiliary head had two DownBlocks (128 and 64 filters), followed by GlobalAveragePooling, batch normalization, a fully connected layer with 64 neurons, and an output layer with two regression outputs. Diffusion step values were encoded with a sine-cosine positional embedding and processed through two fully connected layers with SiLU activation.
We used the Huber loss function (Eq. 4) with $\delta=0.05$ for both tasks, trained with the AdamW optimizer [13], starting with a learning rate of $10^{-5}$ for 256 epochs. The learning rate was halved after 18 epochs without improvement in SSIM (denoising) or $L_{2}$ (regression) on validation.
Two augmentations were applied: MixUp [6] with a probability of 0.25 and a maximum added pattern weight of 0.25, and 10% rate of Gaussian noise-only images paired with zero targets during the denoising phase. $s_{\text{noise}}$ and $p_{\text{noise}}$ were set to 0.5 during the experiments.
Normalization steps were omitted for brevity in Algorithms 2 and 3, but $x_{t}$ values were normalized to zero mean and unit variance. The decoder’s final activation was tanh, scaled by 3.5, while the quality prediction head used a sigmoid activation. $T=384$ in all experiments.
The model was implemented in PyTorch 2.6 with mixed-precision training (bfloat16), and experiments ran on two Nvidia A10 GPUs with 24GB VRAM.
$\mathcal{L}(\hat{x},x)=\begin{cases}0.5(\hat{x}-x)^{2},&\text{if }|\hat{x}-x|\leq\delta,\\ \delta\left(|\hat{x}-x|-0.5\delta\right),&\text{otherwise.}\end{cases}$ (4)
Figure 5: The results of image restoration. N - the number of restoration cycles (the maximum number is $T=384$). Step is the model’s diffusion step estimation (the lower the more restored the image is according to model estimation). Initial quality (estimated quality) is the predicted quality of the original pattern $\hat{q}$; denoising quality - the model estimation of cross-correlation result between restored image and the ground truth $\hat{r}$. The a) and b) - EBSD patterns for Nickel sample of different quality; c) - instrumental noise sample.
5 Results and discussion
In this section, we evaluate the impact of different components of our approach on the denoising performance. Specifically, we assess the contributions of various training and inference strategies used in our diffusion-based denoising framework.
To ensure a fair comparison, all experiments are conducted using the same fixed subset of the dataset, with identical hyperparameters, augmentation strategies, and learning rate schedules. The evaluation is performed on a hold-out test set, where we measure image quality using Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM).
Since our model is trained to predict a denoised image directly from any diffusion step, we first compare different training strategies for the denoising component. This evaluation is based on the model’s ability to directly reconstruct the clean image from a noisy input. Table 1 summarizes the results obtained using different methods for generating noisy training samples, $x_{t}$.
Process PSNR ($\uparrow$) SSIM ($\uparrow$)
Raw data 14.1317 0.5235
Baseline [19] 15.7447 0.6632
Our method (Algorithm 1) 17.6460 0.7728
Table 1: Ablation study of different training procedures for diffusion-based denoising. Raw data represents the test dataset without denoising. Baseline corresponds to the conventional linear algorithm from [19]. Our method follows Algorithm 1, demonstrating significant improvements in both PSNR and SSIM.
The results show that our proposed training strategy significantly outperforms the baseline approach, achieving a noticeable increase in both PSNR and SSIM. This confirms the effectiveness of our training novelties: quality-defined control of training input image generation and linear extrapolation of experimental images towards more degraded state then the experimentally observed image.
Next, we investigate different inference-time denoising strategies, analyzing how modifications to the step selection process impact image quality. Table 2 presents the results of various step selection algorithms.
Process PSNR ($\uparrow$) SSIM ($\uparrow$)
Baseline [19] 17.6737 0.7742
Adjusted ($t_{x}=T\cdot(1-q)$) 17.6771 0.7760
Our method 17.7964 0.7791
Our method (best-step selected) 18.0589 0.7890
Table 2: Ablation study of different denoising procedures. Baseline represents the standard linear denoising approach from [19]. Adjusted modifies the baseline by determining the starting diffusion step based on predicted pattern quality, as $t_{x}=T\cdot(1-\hat{q})$. Our method further improves on this by incorporating dynamic selection of the next diffusion step based on predicted $\hat{r}$, where $t_{\text{next}}=T\cdot(1-\hat{r})$. This method follows the entire Algorithm 3 except for the best-step selection (last line of the Algorithm 3. Our method (best-step selected) refines the previous approach by selecting the final pattern as the one corresponding to the highest predicted $\hat{r}$.
The results indicate that dynamically adjusting the diffusion starting step based on predicted quality leads to consistent performance improvements. The best performance is achieved when the final denoised pattern is selected based on the highest predicted $\hat{r}$, demonstrating the utility of assessing denoising progress in an adaptive manner and is an evidence of the effectiveness of our denoising progress assessment. This result highlights that for some patterns, optimal denoising occurs at an intermediate step, with further iterations potentially degrading quality. The ability to adaptively stop at the best step offers a significant advantage over fixed-step approaches, highlighting the importance of an adaptive stopping criterion for achieving optimal reconstruction while avoiding unnecessary processing.
Figure 6: Predicted values of $\hat{q}$ (pattern quality) and $\hat{r}_{denoise}$ (denoised quality) for a proper dataset, including actual EBSD patterns and images with instrumental noise. The results demonstrate that for real EBSD patterns, the predicted denoising progress values are significantly higher, whereas for noise data, the predicted values remain low for the denoising process.
The observed results in Figure 6 reveal a clear distinction between actual images and instrumental noise samples, based on the predicted values of $\hat{q}$ and $\hat{r}_{denoise}$. For real EBSD patterns, the final predicted denoising progress values are significantly higher, indicating the model’s ability to recognize meaningful patterns. In contrast, for noise data, the predicted values remain consistently low, reflecting the absence of structure in the input data. These findings suggest that the predicted values can serve as a criterion for differentiating between actual images and noise, offering a reliable approach for noise detection and filtering in practical applications.
6 Conclusion
We propose a dynamic method for selecting diffusion steps based on predicted pattern quality, $\hat{q}$, and denoised quality, $\hat{r}_{denoise}$. This adaptive strategy ensures optimal termination of the denoising process, integrates quality-defined control for generating noisy training samples, and extrapolating experimental images to more degraded states, our approach significantly improves denoising performance compared to baseline methods. The predicted values of $\hat{q}$ and $\hat{r}_{denoise}$ also enable noise-only data detection and filtering. These contributions enhance the denoising process, improving reconstruction quality and robustness.
This method could prove useful in future EBSD measurements, allowing for reduced exposure times while sacrificing some pattern quality, and subsequently restoring the information during post-processing. Furthermore, the suggested approach lessens the reliance on the experimentalist’s selection of the image processing sequence and could be useful beyond EBSD context.
7 Acknowledgments
We would like to acknowledge the partial support of our computational resources through the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-9949.
Collecting EBSD patterns is performed at AMOLF, an NWO funded institute, and is part of the project ”Achieving Semiconductor Stability From The Ground Up” (with project number 19459) which is financed by the Dutch Research Council (NWO), Gatan (EDAX), Amsterdam Scientific Instruments (ASI) and CL Solutions.